The mean salary offered to students who are graduating from Coastal State University this year is $ 24,260 , with a standard deviation of $ 3643 . A random sample of 80 Coastal State students graduating this year has been selected. What is the probability that the mean salary offer for these 80 students is $ 24,500 or less? Carry your intermediate computations to at least four decimal places. Round your answer to at least three decimal places.
Answer #1

Let X be the salary offered to students who are graduating from coastal state university.

Given:

u= 24260

0=3643

n=80

The probability that the mean salary offer for these 80 students is $24500 or less given by

24500 - u PT < 24500)=PI-

PT < 24500) = P(Z < 24500 - 24260 3613 80

240 Pī < 24500) = P(Z < 8.9443

240 Plī < 24500) = P(Z <-

Plī< 24500) = P(Z < 0.5892)

Plī< 24500) = P(Z < 0.59)

Plī< 24500) = 0.7224

### by using z table

Hence required probability is:

Plī< 24500) = 0.7224